You’ve observed the following returns on Yamauchi Corporation’s stock over the past five years: –29.7 percent, 16.8 percent, 36.6 percent, 3.9 percent, and 22.9 percent. a. What was the arithmetic average return on the stock over this five-year period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.) b. What was the variance of the returns over this period? (Do not round intermediate calculations and round your answer to 6 decimal places, e.g., 32.161616.) c. What was the standard deviation of the returns over this period? (Do not round intermediate calculations and enter your answer as a percent rounded to 2 decimal places, e.g., 32.16.)
Answer #1
Year Return % Deviation from mean [d] d^2
1 29.7 7.72 59.5984
2 16.8 -5.18 26.8324
3 36.6 14.62 213.7444
4 3.9 -18.08 326.8864
5 22.9 0.92 0.8464
109.9 627.9080
a] Arithmetic average = 109.9/5 = 21.98%
b] Variance = 627.9080/5 = 125.581600
c] Standard deviation = (125.5816)^0.5 = 11.21%